OpenStudy (anonymous):

Differentiation Question:

OpenStudy (anonymous):

According to Wolfram Alpha the derivative of \[\sin ^x(x)\] is \[\sin ^x(x)(x \cot (x) + \ln (\sin (x)))\] but when I work it out I get \[\ln (\sin (x))+\cot (x)\ln (x)\] Could someone explain if how they're equal, if they are actually equal?

OpenStudy (anonymous):

@Hero ? @Preetha ?

OpenStudy (anonymous):

they are not equal

OpenStudy (anonymous):

\[\sin^x(x)=e^{x\ln(\sin(x))}\] take the derivative of that using the chain rule and the product rule

OpenStudy (anonymous):

is that what you did, or did you "take the log" you get the same answer either way, all the work is taking the derivative of \[x\ln(\sin(x))\]

OpenStudy (anonymous):

@satellite73 Ok, thanks, I see my error now.